OpenStudy (mathgenus123):

plz help i will medal and fan Bill says the solution to 3(x – 1) +6 = 12 and 3x + 3 = 12 are the same. * A. Is Bill correct? B. Show your work to justify your answer.

OpenStudy (anonymous):

If they have the same solution then wouldn't they be equal to each other?

OpenStudy (anonymous):

Do you know how to factor

OpenStudy (anonymous):

So does 3(x-1)+6=3x+3

OpenStudy (anonymous):

I think mathgenus123 wants to know how to solve the first equation

OpenStudy (mathgenus123):

Simplify the expression by Combining Like Terms 4x + 2x + 4x + 4Y – Y + 2x *

OpenStudy (anonymous):

By factoring the 3 times x you get 3x and 3 times -1 is -1 so you have 3x-1+6=12, see if you can solve from there if not I can help

OpenStudy (anonymous):

actually \[3\times(-1)=-3\] so you would have 3x-3+6=12